Solve: y – 4y + 3y = 9t – 3 y(0) = 3, y(0) = 13 y(t) = Preview
Answer #1

Given DE, y-4y+3y=9t-3° 4 (0) = 3, y 10) = 13. Apply LoT on both 0 sides. hy - ugl+34} = (49t-3} L241} -4224} + 3 L2 y} = 9353+52-93+9 - A + * + Šat 5-3 st szt + D S2 5² (34) (5-3) 353 +5²-gst s cs_12153) TOSES) As. (S-1) (5-3) & B (S4) (-3) + C CS10 =-26 > C= -185 » [[= -5] put S= 3 in a ③ we get vale for D 3 (3)3 + (3) 2 _ 3 (8) +9 = to 8o (CO) to (3² (3-1) 81+94 +9comment me have dont any doubt forget for please rating